### v2

# Uncertainty-driven Adaptive Exploration

Abstract
Adaptive exploration methods propose ways to learn complex policies via alternating between exploration and exploitation. An important question for such methods is to determine the appropriate moment to switch between exploration and exploitation and vice versa. This is critical in domains that require the learning of long and complex sequences of actions. In this work, we present a generic adaptive exploration framework that employs uncertainty to address this important issue in a principled manner. Our framework includes previous adaptive exploration approaches as special cases. Moreover, we can incorporate in our framework any uncertainty-measuring mechanism of choice, for instance mechanisms used in intrinsic motivation or epistemic uncertainty-based exploration methods. We experimentally demonstrate that our framework gives rise to adaptive exploration strategies that outperform standard ones across several MuJoCo environments.
1 Introduction
Over the past decade, deep reinforcement learning (DRL) [14, 19, 32] has become the policy learning method of choice in a variety of domains. In online DRL, the goal is to approximate the $Q$-function via alternating between generating new experiences by interacting with the environment (exploration) and updating the strategy using this experience (learning) [23].
Now, robotics is considered a standard domain for testing DRL [10, 13, 14, 18]. In general, such domains are characterized by: (i) continuous and multi-dimensional action space; (ii) informative, multi-component, and dense reward function (iii) episodes that are very probable to be terminated early (or “truncated”) when an agent is set in an unsafe state; (iv) transitions being highly deterministic. In such domains, the main aim is to learn a complex and long trajectory, necessitating the execution of a sequence of actions that will lead the agent to a specific state.
In domains with the aforementioned characteristics, it is critical for an agent to pose the question: “Is it always better to search for new states, or should I carefully choose when to explore?” To further understand the importance of correctly answering that question, let us consider the classic Walker domain [31]. In this domain, a two-dimensional bipedal robot aims to walk in the forward direction. At the beginning of an episode, the agent is set to an initial state. An episode is truncated when the agent finds itself in an unsafe state. Suppose also that the agent has already learned an incomplete and close-to-optimal trajectory which, in this case, corresponds to performing a single step. This trajectory can guide the agent from the initial state to a new safe one. The agent’s goal is to extend this trajectory as much as possible. For that to happen, two phases are required: (i) the exploitation of the current policy by replicating the trajectory until its end (by replicating in other words the “single step” trajectory); and (ii) the exploration from the end of the trajectory until finding the next best action (explore after performing a “single step”). We should note that the switch between the two phases should happen within each episode. In case the agent explores instead of replicating the known trajectory, it is more probable to transition to a worse—and maybe unsafe—state rather than the one included in the trajectory. Given this example,  Pislar et al. [23] argue that it is more advantageous to ask “when to explore” rather than “how to explore”. [^1] The latter question is in fact the question that deep exploration methods [3, 15, 17, 21, 25] focus on. Even though these were designed for use in structurally different domains (e.g., video game domains with sparse, discrete rewards, and stochastic transitions), they are regularly used in robotics as well ([17, 18]).
In this paper, we put forward ADEU - ADaptive Exploration via Uncertainty, a generic exploration framework that is able to decide in a principled manner when the switching between exploration and exploitation and vice versa should take place within an episode. Intuitively, ADEU enables exploration when it adjudges that its uncertainty regarding its policy is “high enough”. Now, uncertainty is arguably a subjective term and is highly dependent on the formulation of the problem being solved. ADEU takes that into account. It can in fact incorporate any mechanism for measuring uncertainty in its action selection mechanism. Intuitively, such an uncertainty-measuring mechanism corresponds to a particular “uncertainty definition”. Moreover, ADEU does not resort to the use of heuristic techniques or thresholds to decide the aforementioned “is uncertainty high enough” question: instead, it selects an action in a principled manner, via sampling from a distribution whose spread is proportional to its uncertainty regarding the policy. By being able to incorporate any uncertainty-measuring mechanism, and by selecting an action in the aforementioned principled manner, ADEU is in fact able to generalize over existing adaptive exploration strategies.
To test our framework, we create $3$ different instances of ADEU, each one corresponding to $3$ different uncertainty-measuring mechanisms. We test our method in the most complex/difficult among the robotic tasks in [31]. Our experiments demonstrate that the performance of an ADEU instance depends on the ability of each uncertainty-measuring mechanism to quantify uncertainty. However, in our experiments, the instance of ADEU employing a specific uncertainty-measuring mechanism, outperforms the original exploration method that employs the corresponding uncertainty-measuring mechanism. Moreover, there is always some instance of ADEU that performs better than all of its opponents. As such, ADEU can be considered as a ‘plug-and-play’ adaptive exploration framework, that can be coupled with any uncertainty-measuring mechanism; and exhibits the ability to improve the performance of DRL algorithms that employ the corresponding uncertainty-measuring mechanism in their exploration process.
2 The ADEU framework
A Motivating Example:
Let us consider a variation of the classic Frozen Lake example as shown in Figure 1.222The official documentation of the Frozen Lake environment can be found at https://gymnasium.farama.org/environments/toy_text/frozen_lake In that environment, the agent has to discover the sole trajectory that will eventually lead it to the goal passing through frozen tiles, while if it steps on a “hole” tile the episode terminates. In our example, the agent is rewarded when it reaches the next frozen tile. This reward is inversely proportional to the distance of the goal: the reward function outputs a higher reward when the agent reaches an allowed tile that is also closer to the goal. In addition, let us assume that an agent has already been trained to reach state $61$ (state at the end of the red line in Figure 1—right above state $71$). In other words, its policy $\pi(s)$ can successfully lead it to state $61$, replicating the red trajectory. The agent aims to extend the red trajectory until reaching the treasure.
Figure 1: The Frozen Lake environment, in which there is a single path connecting the initial state with the target. Exploration becomes more challenging when increasing the grid dimensions.
An effective way for that agent to explore this environment is by “trusting” its policy until the state $61$ and then switching to exploration. Assume an agent employing a classic exploration technique like $\epsilon$-greedy (or any noise-based strategy) that explores every episode with a constant $\epsilon$ (or constant variance in the case of noise-based strategy). By not modifying $\epsilon$ during an episode, the agent explores with a constant rate. Then, the agent will probably not be able to replicate the red trajectory since it is probable (approximately $\epsilon$% at each timestep) that the agent will deviate from the known policy.
In an even worse scenario, assume an exploration strategy that forces the agent to visit states with high uncertainty (e.g., a state that has not been visited enough times, or a state whose value the agent is unsure about). In that scenario, the agent is uncertain for all the states not belonging to the trajectory. Thus, this exploration strategy will force the agent to needlessly explore rather than trusting the trajectory, leading to premature termination of large number of episodes before managing to actually extend the trajectory. The use of a greedy-in-the-limit exploration strategy would lead to similar “over-exploration” phenomena.
Now, suppose an adaptive exploration method that determines at every step whether (i) the agent should explore a state, or (ii) it should exploit/trust the known policy for that state. By appropriately alternating between these two phases, the agent can replicate the known trajectory and subsequently begin exploration. A key challenge lies in determining how the agent identifies the appropriate moment to transition between the phases.
Let us define $f(s)=\frac{1}{\beta\sqrt{n(s)}}$ where $n(s)$ is a (pseudo-)counter that counts (or approximates in larger environments) the times that a state $s$ has already been visited and $\beta<1$ a positive constant. Let us also suppose that the states of the trajectory have been visited numerous times. Then, $f(s)$ arguably provides an effective signal on whether to switch (or alternate) between the two phases. This indicates that an agent should explore proportionally to the value of $f(s)$.
2.1 Formal Framework Description
Against this background, we put forward ADEU, a framework allowing for adaptive exploration, based on the uncertainty associated with each state. ADEU makes use of some uncertainty-measuring mechanism $f(s)$, whose value dictates when an agent should start exploring and when it should trust the gained knowledge, similarly to the example. As discussed earlier, uncertainty is a subjective term; its definition is closely related to the problem being solved. The ADEU framework implicitly accepts any definition of uncertainty via the implementation of an $f(s)$. As such, practically all uncertainty-measuring mechanisms—e.g., those utilized by intrinsic motivation to detect novel states, or by epistemic uncertainty-driven exploration to quantify the convergence degree of a policy—are compatible with ADEU. We will elaborate on this after formally presenting the key ADEU principles.
The ADEU framework can be summarized concretely by Equation 1, which defines the way that an ADEU agent selects an action given a state $s$.
$a(s)\sim\mathcal{D}\left(\pi(s),g(f(s))\right)$ (1)
In more detail, this action will be chosen by sampling from a distribution $D$. The first moment of that distribution is the policy $\pi(s)$ that the DRL agent has already learned. Hence, the mean of the probability distribution $D$ (intuitively, its center of mass) is the policy $\pi(s)$. The second moment (variance) of that distribution is $g(f(s))$, which is proportional to the uncertainty. Hence, $g(f(s))$ defines how “spread” the distribution is, and how close the sampled action will be to the policy $\pi(s)$.
The key intuition in ADEU is that the spread of the distribution should be defined so that it corresponds to the agent’s uncertainty regarding the $\pi(s)$ policy; while that uncertainty can be calculated in any way considered appropriate. [^3] In this paper, we use the first raw moment (mean) and the second central moment (variance), both in the context of continuous and discrete distributions; but one could conceivably use any other moments of choice to enhance the action selection. For instance, one could use some function of the third standardized moment in order to guide exploration towards a specific direction, e.g., a direction that is deemed safe [33, 34], if such a direction can be in some way associated with the kurtosis of the distribution. To this end, $f(s)$ in $g(f(s))$ is some mechanism able to compute uncertainty, like the ones used in uncertainty-based exploration methods[4, 13, 21] or mechanisms designed for intrinsic reward construction in intrinsic motivation methods[1, 3, 17]. [^4] To clarify, in our method, we use the mechanism used in intrinsic motivation methods for defining $f(s)$, and not for providing an exploration bonus. By default, $f(s)$ will be outputting a positive scalar. In multidimensional distributions $D$, this scalar can multiply an Identity matrix. In addition, since $f(s)$ can be any uncertainty-measuring mechanism, we use $g(\cdot)$ as a normalizer to prevent $f(s)$ from outputting enormous values.
Moreover, $D$ can be any distribution deemed appropriate. In this paper, we use the Multinomial distribution in discrete action space experiments/examples, and the Gaussian (or Normal) distribution in continuous action space ones. We selected these distributions due to their simple implementation.
Guided by Equation 1 to select actions, an ADEU agent is able to alternate between exploitation and exploration within the episode in a principled manner, and its behaviour can be summarized as follows. At every timestep, the agent (i) calculates the uncertainty $f(s)$ regarding that state; (ii) constructs a distribution $D$ in which the first moment is the current policy $\pi(s)$, and the second moment is the $g(f(s))$; and (iii) samples from that distribution. In case of low uncertainty, the distribution will not be “spread” and hence the sampled action will be close to $\pi(s)$. Notice that this effectively causes the agent to replicate parts of a trajectory corresponding to a $\pi$ it is confident about. In any other case, the sampled action will probably be far from $\pi(s)$, leading the agent to explore new actions in that state. This new experience can potentially lead to an extended trajectory.
Note that the discussion above does not imply that ADEU needs to be provided with any trajectories or other forms of background knowledge. ADEU assumes a zero-length initial trajectory at the beginning of the training process, and then is progressively able to extend it via its adaptive exploration technique. [^5] Of course, if ADEU is provided with an accurate initial trajectory or policy, it can exploit such knowledge. In other words, at the start of training, the agent has not yet converged at a policy, even for the initial states. As such, it lacks certainty about which actions to take and must explore until it converges to a reliable policy even for the initial state. Once this occurs, the agent follows the learnt action policy from the initial state until it reaches the next state where “enough” uncertainty arises. There, it resumes exploration. This cycle of exploration and policy refinement continues until the agent becomes confident across the entire trajectory.
ADEU’s efficiency arises from the fact that it does not needlessly explore, but instead replicates the parts of a “known” or “learnt” trajectory that it trusts. But this trajectory might be sub-optimal. So, to keep improving the learnt trajectory, ADEU also performs the so-called rollout episodes. In more detail, at the beginning of every episode, with probability $\rho$ ADEU performs exploration using as uncertainty mechanism the $f_{\text{rollout}}(s)=c$, where $c$ is a constant number. Hence, for that specific “rollout” episode, ADEU takes the chance to explore so as to improve the learnt trajectory. We should note that in the case of a well-constructed $f(s)$ that smoothly decreases, $\rho$ should and can simply be set to 0. For simplicity, in our implementation we treat $\rho$ as a constant that does not change throughout the training process. Also, we set $f_{\text{rollout}}(s)=c$ to present a simple version of our framework, but, in more complex environments, this can change too.
2.2 Generality of the ADEU framework
We now elaborate on the fact that ADEU is a generic adaptive exploration framework. First, it can readily use as $f(s)$ any predefined mechanism—e.g., some epistemic uncertainty-measuring mechanism, or mechanisms used in intrinsic motivation methods to approximate visitation frequency. Moreover, it is a framework that can emulate most, if not all, existing adaptive exploration techniques. This can be achieved by designing an as-complex-as-required $f(s,\cdot)$— e.g., an $f(s,\cdot)$ mechanism that outputs a value considering previous states or actions, or other constraints.
Using predefined mechanisms as $f(s)$:
To see this, consider first, the intrinsic motivation (IM) family, which rewards the agent whenever a novel state is visited. This non-stationary reward, termed as $r_{\text{intrinsic}}(s)$, is added to the external reward provided by the environment, as described in Equation 2. Then, the agent acts greedily to maximize 2.
$r(s)=r_{\text{extrinsic}}(s)+r_{\text{intrinsic}}(s)$ (2)
To calculate the $r_{\text{intrinsic}}(s)$, IM-oriented methods [1, 3, 17, 18] propose numerous mechanisms that can quantify the novelty of a state. For example, RND [3] proposes to calculate the visitation frequency as the approximation error between two networks named Target and Predictor. Predictor network is trained to predict the Target’s output for a given state. The difference between the output of the two networks (Equation 3) approximates the visitation frequency for that specific state.
$r^{\text{RND}}_{\text{intrinsic}}(s)=\left\lVert f_{\text{predictor}}(s)-f_{\text{target}}(s)\right\rVert_{2}^{2}$ (3)
ADEU can utilize this approximation error as an uncertainty-measuring mechanism. That is, an instance of ADEU that defines uncertainty as the visitation frequency of a state can utilize $f(s)=r^{\text{RND}}_{\text{intrinsic}}(s)$. Of course, any other mechanism approximating visitation frequency, such as the ones proposed in other IM methods [1, 17, 26] can also be employed.
One can also create an instance of ADEU that utilizes an uncertainty mechanism similar to the ones defined in epistemic-uncertainty driven methods [4, 13]. Epistemic uncertainty-driven methods employ different representations of the same Q-function to measure how certain an agent is about its policy in a specific state. For example, the UCB [4, 13] strategy greedily selects actions that maximize both the $Q_{\text{mean}}$ and the $\lambda Q_{\text{std}}(s,a)$ as shown in Equation 4.
$a=\arg\max_{a}\left[Q_{\text{mean}}(s,a)+\lambda Q_{\text{std}}(s,a)\right]$ (4)
where $Q_{\text{mean}}(s,a)$ is the mean $Q$-value calculated across the different representations of the $Q$-function, $Q_{\text{std}}(s,a)$ is the corresponding standard deviation, and $\lambda$ is a positive constant. Given this, we can also create an ADEU instance that mirrors the above uncertainty-measuring method, via setting $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$. Table 1 summarizes the discussion above.
Table 1: Instances of ADEU using different uncertainty-measuring mechanisms—examples only; any IM mechanism and any mechanism calculating confidence over $\pi$ can be used.
Definition of Uncertainty An example of $f(s)$
Uncertainty is calculated regarding the novelty of each state $f_{\text{RND}}(s)=\left\lVert f_{\text{predictor}}(s)-f_{\text{target}}(s)\right\rVert_{2}^{2}$
Uncertainty is calculated as the confidence of an agent about its policy $f_{\text{UCB}}(s)=\lambda Q_{\text{std}}(s,\pi(s))$
Emulating adaptive exploration techniques:
Table 2 presents certain key adaptive exploration works, and their instantiation within the ADEU framework. See also Section 4 for more details.
Table 2: Previous adaptive exploration works and how these can be instantiated in ADEU.
Previous Work Description Special Case of ADEU
Value-Difference Based Exploration[29] Uses $\epsilon$-greedy. Modifies $\epsilon$ at every state wrt. TD-Error A Bernoulli distr. can be used to simulate $\epsilon$-greedy. f(s)f(s)=TD-Error(s)
$\epsilon z$-greedy [6] Agent performs an “option” (i.e, a sequence of actions) with constant probability $\epsilon$. The length of that option is sampled from a heavy-tailed distribution. With prob. $1-\epsilon$, agent uses its policy $\pi(s)$. We will use a Bernoulli to simulate whether an agent will act according to its $\pi(s)$ or not. Then $f(s,n)$, where $n$ is sampled from the heavy-tailed distribution at the first call of $f(s,)$, will change Bernoulli’s pp. [^6] At the first call, $f(s,n)$ where $n$ is an optional argument not used in the first call, it will sample the $n$. Then, at every next call of the $f(s,n)$ the $n$ parameter will be used as an argument and decreased by $1$.
Pislar et al. [23] This is a general work discussing ‘when should agents explore’. It proposes intra-episodic exploration after an agent receives a signal. This exploration ‘session’ lasts for a fixed number of timesteps. A Bernoulli is required to alternate (by modifying $p$) between $\pi(s)$ and a random action. $p$ will be modified according to $f(s)$, which will replace the signals. After $f(s)$ changes $p$ to promote exploration, it will output a constant value for a fixed number of timesteps.
Goal-based exploration (“Go-explore” approach) [8, 9, 30] An agent first replicates a trajectory (or a ‘return’ or ‘go’ phase), and then explores. Either a Bernoulli or a Gaussian can be used, depending on the action space. $f(s)$ will be a binary function.
Table 2 explains how ADEU incorporates previous works on adaptive exploration. Given that ADEU can subsume existing adaptive exploration methods as special cases, it is reasonable to expect that it can replicate their performance in the tested domains. Thus, ADEU is well-suited for tasks with both dense and sparse reward structures. [^7] Regardless, we plan to test ADEU in environments with sparse reward functions also, like ATARI2600 [2]. In addition, by using more complex definitions of “uncertainty” (e.g., Nikolov et al. [20], Badia et al. [1]) we can enhance exploration in domains with less informative and sparse reward functions.
Motivation example cont’d:
We now continue with the technical setup and the results of the motivation example. We design a large-scale grid environment of size $1500\times 1500$, extending the setup shown in Figure 1. In this environment, we evaluate four different agents: (i) a standard $Q$-learning agent using $\epsilon$-greedy exploration, (ii) an intrinsic motivation (IM) agent built on top of $Q$-learning, (iii) a UCB-based ensemble agent composed of multiple $Q$-learners, and (iv) an instance of ADEU. For the IM agent, we adopt ideas from [26] without the need for approximation due to the discrete state space. The ADEU agent adjusts its action selection based on state visitation frequency, favoring exploration in less-visited states and exploiting learned values in frequently visited ones. It uses the same mechanism as the IM agent. Additional implementation details about the environment, the probabilistic distribution and activation function choices are provided in the Appendix.
Table 3 shows the maximum reward and the mean reward during learning, computed as follows. We ran $5$ different runs (seeds) for all agents, and the following process was performed for each of the runs: Once every $100$ episodes, an “evaluation” episode was executed, in which the agent was using greedily its policy (i.e., no actions were selected for exploration during that episode); and we stored the reward accumulated during that episode, and all such episodes in each run. We then create an “average run” by averaging over these episodes’ values across all runs; and compute the mean episodic reward and the maximum episodic reward in that average run. Note that the maximum episodic reward was achieved after convergence.
Table 3: Mean and Max episodic rewards of different exploration strategies in the Frozen Lake domain. Results are averages over 55 runs. ADEU was built on top of a $Q$-learning agent.
Exploration Strategy Mean and Max episodic rewards
$\epsilon$-greedy 334.0; 983.0
ADEU (Using visitation frequency) 88,854.0; 104,487.0
IM 393.0; 989.0
UCB 36,070.0; 104,487.0
As shown in Table 3, ADEU demonstrates superior performance compared to the other methods by rapidly learning a policy that consistently guides it toward the target price. Specifically, the ADEU agent avoids excessive exploration of frequently visited states—those characterized by low uncertainty according to the employed function $f(s)$. Instead, it tends to replicate the learnt trajectory until completion before initiating further exploration; and manages to learn the correct trajectory. By contrast, both the $\epsilon$-greedy and IM algorithms fail to learn a successful trajectory to the goal, likely due to their persistent over-explorative behavior. Finally, UCB learns the correct trajectory, but it requires nearly four times as many episodes as ADEU (16,800 vs 4,515 for ADEU). Its average reward is approximately half that of ADEU’s. Note that in our main experiments involving more complex action and state spaces, UCB always ranks behind at least one ADEU variant.
Figure 2: (Top) Blue line shows $\pi(s)$ across one episode. The red line shows the action selected by the agent. (Bottom) Uncertainty as calculated by the agent across the states of a single episode.
Figure 2 (top figure) shows both the actual policy $\pi(s)$ and the action selected from the ADEU agent (so, using Equation 1) during a single episode in our Frozen Lake environment. In addition, the bottom figure shows the value of $g(f(s))$ during the same episode. As expected, at the beginning of the episode $g(f(s))$ outputs low values since the agent visits states that it is certain about. Hence, the action selected by the ADEU agent coincides with the one dictated by $\pi(s)$. As the agent visits states with high uncertainty, it does not trust its so far learned policy and starts exploring them. That is, the action selected by ADEU and the policy $\pi(s)$ differ at the end of the episode, since $g(f(s))$ increases. We should also note that the action suggested by $\pi(s)$ in the later timesteps of the episode is not the ideal one. However, the action sampled for the exploration strategy was the correct one, allowing the agent to “survive” for more timesteps and gain useful experience. Thus, indeed, the exploration phase should have begun at the later timesteps; had it started earlier, the agent would not have “survived” to extend the known optimal policy for the early timesteps to the later ones.
ADEU’s variants for safe exploration
Now, as discussed and seen in the example, ADEU avoids exploration in states with low uncertainty. Moreover, it can be extended so that it conducts safe exploration—i.e., to not explore in unsafe states [33, 34]. By simply defining a new mechanism that also considers this extra constraint, ADEU can avoid exploring these states.
More precisely, suppose that a mentor has already informed the agent about some potential unsafe states (like in Thomas et al. [27]) and the $\pi(s)$ policy the agent should follow in these particular states. We would like the agent to not explore finding a better policy in these specific states. Hence, by designing an $f(s)$ that outputs a low uncertainty in these states, an ADEU agent can alternate between exploration and exploitation proportional to the safety of the current state.
Similarly, the agent can be informed of the cost function of the environment (like in Tian et al. [28]). By augmenting $f(s)$ with the cost function, ADEU can combine adaptive exploration with safety. Even if the ADEU agent does not obtain any knowledge regarding the unsafe states or the cost function, it can interact with the environment to approximate it. Along with this “approximation” process, the agent can still respect the safety constraints.
3 Experimental Setup and Results
To experimentally test our framework, we create two main instances of ADEU, using as $f(s)$ (i) the intrinsic motivation mechanism proposed in [3] and (ii) the epistemic uncertainty measuring mechanism proposed in [4]. These are the agents described in Table 1. Both agents were built on top of a TD3 [11] DRL algorithm’s original implementation. [^8] https://github.com/sfujim/TD3. As such, we refer to them as: (i) TD3+ADEU with f(s)=RNDf(s)=RND; [^9] More precisely but less concisely, “TD3+ADEU with $f(s)=r_{\text{intrinsic}}^{RND}$”. and (ii) TD3+ADEU with $f(s)=Q_{\text{std}}\left(s,\pi(s)\right)$.
In addition, we consider an instance of ADEU with a non-informative $f(s)$, and create a ‘TD3 + ADEU with $f(s)=c$’ agent, where $c$ is a constant. Obviously, since $c$ is a constant, $f(s)=c$ is not an appropriate uncertainty measuring mechanism. The agent with constant $f(s)$ will serve as baseline since its distribution matches the one used in TD3 [11] and hence, they share the same exploration strategy. We use $g(f(s))=c$ for this ADEU instance. For the other ADEU instances, we use Equation 5, where $c$ has the same value as in the naive ADEU instance. That value was set to $0.2$, as in the TD3 implementation.
$g(f(s))=sigmoid(f(s))\cdot c$ (5)
As mentioned, our agents use TD3 as a “basis” algorithm. In principle, any DRL algorithm could have been used instead (e.g., any DRL method well-suited for continuous action spaces, if the focus is in such domains). Since ours is an exploration framework, we wanted to use a DRL algorithm that is not tied to any exploration strategy—and TD3 uses a ‘vanilla’ one that can be readily replaced with ADEU. By contrast, SAC [12] uses the entropy framework, and hence it is tied to that framework’s particular exploration strategy.
We also implement three agents to act as our main competitors, in order to test ADEU against representatives of three standard families of exploration methods: IM exploration, epistemic uncertainty-driven exploration, and (parameter-space) noise exploration.
Specifically, we test against TD3+RND, an agent that acts greedily in the environment with a reward function equal to $r_{\text{extrinsic}}(s)+r^{\text{RND}}_{\text{intrinsic}}(s)$ (cf. Equations 2 and 3). In addition, we compare against TD3+UCB, an agent that selects actions aiming to maximize Equation 4.
Finally, for interest, we also create an agent that uses Noisy Nets [10] to perturb the actor’s network and achieve more efficient exploration. We compare against this TD3+Noisy Nets agent, since Noisy Nets is a popular choice for exploring robotic domains. Note that Noisy Nets does not utilize any definition of uncertainty—it is just an “improved” TD3 agent in terms of exploration.
We conduct our experiments in standard MuJoCo domains, provided by [31]. Specifically, we used the five domains considered as the most difficult ones according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ In these domains, the agent needs to explore in the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., [1, 20]. Note that we did not use [7] since their domains do not have the early termination condition.
Domain TD3 + ADEU TD3+UCB TD3+RND TD3+Noisy Nets
$f(s)=c$ $f(s)=\text{RND}$ $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$
Walker2d 564; 726 771; 980 411; 488 366; 410 5; 28 489; 631
Hopper 412; 624 716; 1127 301; 568 226; 425 13; 44 495; 912
Swimmer 40; 46 52; 62 36; 39 32; 36 -3; -2 30; 35
Ant 345; 980 1072; 1341 1134; 1500 585; 790 -2658; 980 364; 981
Humanoid 86; 132 313; 345 512; 658 89; 149 85; 132 75; 88
Hum. Stand. 61511; 64490 67881; 73303 80277; 97208 60878; 62227 47783; 47928 62152; 63645
Table 4: Mean and Max episodic rewards in various domains [31]. Results are averages over $15$ runs.
Table 4 presents agents’ performance in terms of mean and max episodic accumulated reward in the aforementioned domains. These mean and max values were calculated over $15$ runs (seeds), via the process described in Section 2.2. We mark in bold the best-performing algorithms in terms of mean episodic reward within each domain—an algorithm is considered to be “best-performing” if its mean episodic reward reaches at least $95\%$ of the highest mean episodic reward accumulated in the domain.
ADEU improves DRL performance:
As seen in Table 4, TD3+ADEU with $f(s)=RND$ or TD3+ADEU $f(s)=\lambda Q_{std}(s,\pi(s))$ outperform the others in every domain. In addition, in the three most difficult domains (i.e., Ant, Humanoid Standup, and Humanoid) both of those ADEU instances outperform all other agents, achieving the first and the second best performance. It is also noteworthy that an instance of ADEU equipped with a specific uncertainty-measuring mechanism $f(s)$, typically outperforms the original exploration method that employs the corresponding uncertainty-measuring mechanism.
In addition, our results show that ADEU combined with $\lambda Q_{\text{std}}(s,\pi(s))$ outperforms most, if not all, of its opponents in the highly challenging Ant, Humanoid, and Humanoid Standup environments. That indicates that this mechanism tracks uncertainty more efficiently in larger environments.
ADEU is a ‘plug-and-play’ adaptive exploration framework:
As seen in our experiments, ADEU can incorporate any uncertainty-measuring mechanism of choice. Of course, as also discussed above, not all uncertainty-measuring mechanisms are created equal. Notice, e.g., that the naive instance of TD3+ADEU, TD3+ADEU with $f(s)=c$, is outperformed by at least one other ADEU instance in each domain. This indicates that the efficiency of ADEU depends on the efficiency of the uncertainty mechanism. Seen otherwise, the choice of an appropriate uncertainty-measuring mechanism for use in the ADEU framework, promises the improvement of performance in the domain of interest.
4 Related Work
Intrinsic Motivation: Several IM approaches have been tested in robotic domains. For instance, Lobel et al. [17] employ a Coin Flip Network to design an efficient intrinsic motivation function that can be utilized as an uncertainty mechanism in ADEU. They test their method in a sparse reward variance of the Fetch domain, which is more comparable to deep exploration settings. On the other hand, [18] adds a random reward to the sparse reward function of the environment, to guide the agent to explore different parts of the environment. This can be used for testing our framework in sparse reward function domains.
Epistemic Uncertainty: As mentioned earlier, UCB [4, 13] is commonly used in continuous action-space robotic domains. Inspired by these works, we also created an instance of ADEU using this ensemble mechanism to track uncertainty (the ADEU instance with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ shown in Section 3) and test it against a UCB implementation built on top of a TD3 agent. Similarly, [5] measures uncertainty as the variance of the different $Q$ representations. In the case of high uncertainty, the ensemble agent “asks” a mentor for guidance.
Adaptive exploration: Tokic et., al [29], uses the Temporal-Difference Error to change the $\epsilon$ parameter of the $\epsilon$-greedy in each state. Hence, an agent is more likely to explore when uncertainty about the $Q$-function is high. Pislar et., al [23] is the first work that actively attempts to create a modern framework for intra-episodic exploration. They argue that an agent should switch from exploitation to exploration when an external signal arrives. They propose numerous “signals” for changing modes, either heuristic or random ones. As noted earlier, these signals can be used as uncertainty mechanisms in ADEU. Similarly to [23], ez-greedy [6] suggests following directed behavior with random—sampled from a heavy-tailed distribution—length. This directed behavior, guides the agent to repeat a single action numerous times. Thus, it achieved excellent results in settings which require replicating the same action numerous times, such as DeepSea [22] and Chain Walk [21]. On a similar note, [16] proposes an adaptive exploration variant of PPO [24], where the entropy term in the loss function is scaled by a heuristic based on the agent’s recently obtained rewards. Even though ADEU does not directly manipulate the loss, it can use this heuristic as $f(s)$. Finally, most goal-based exploration works [8, 9, 30] can be considered as ‘adaptive exploration’ ones since they replicate the known trajectory with no exploration, and then use some exploration to extend it.
5 Conclusions and Future Work
In this paper, we put forward a generic adaptive exploration framework, that uses a generic uncertainty-based action selection mechanism to decide in a principled manner when to alternate between exploration and exploitation and vice versa. The generality of that mechanism allows (i) our framework to effectively incorporate any uncertainty-measuring mechanism of choice; and (ii) existing adaptive exploration techniques to be viewed as special cases of our framework. In problems requiring adaptive exploration, ADEU is a promising and easily deployable solution. Additionally, it allows the user to select an existing uncertainty measurement mechanism or to define a heuristic one tailored to its problem. Our experiments verified that ADEU instances outperform other exploration strategies in standard difficult robotic testbeds.
Ongoing and future work includes performing more tests and extending ADEU in various directions. To begin, we intend to perform tests to further verify the ability of ADEU to exploit background knowledge; and its ability to recover from being fed with sub-optimal policies and erroneous background information (e.g., via the use of the rollout episodes or via designing mechanisms that account for a policy’s value). Finally, we intend to design mechanisms to appropriately set the third and fourth moments of $D$, to guide exploration to particular directions—a useful feature for safe exploration.
Acknowledgments and Disclosure of Funding
The research described in this paper was carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: DEEP-REBAYES. HFRI Project Number 15430).